OpenStudy (anonymous):

Consider the equation 5x + 2y = 30. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

Hold on, Hun. :)

OpenStudy (anonymous):

. The given linear equation is: 5x+2y=30 To find x - intercept we substitute x=0 . 5x+2y 5⋅0+2y 2y y =30 =30 =30 =15 So, the point (0,15) is on the line. To find x - intercept we substitute y=0 . 5x+2y 5x+2⋅0 5x x =30 =30 =30 =6 So, the point (6,0) is on the line. The graph of the function is:

OpenStudy (anonymous):

(0,15) and (6,0)

OpenStudy (anonymous):

Thank you